# AVerImaTeC: A Dataset for Automatic Verification of Image-Text Claims with Evidence from the Web

### v1

@@ 1 Introduction
We further introduce a baseline for image-text claim verification, which operates by generating evidence-seeking questions aimed at fact-checking and answering these questions with a set of expert tools. Since AVerImaTeC is the first image-text claim verification dataset to incorporate QA annotations that explicitly reflect reasoning paths and evidence, we develop a reference-based evaluation method to assess models’ generated questions and retrieved evidence. Using this evidence evaluation, we report conditional verdict accuracy which measures the correctness of predicted verdicts only when the associated evidence score exceeds a predefined threshold.
@@ 2 Related Works
Table 1: Comparison of fact-checking datasets. Indep. (independence) denotes whether extracted claims are context independent (e.g., understandable without fact-checking articles). Img., Suff. and Retr. are abbreviations for image, sufficiency and retrieval, respectively. Sufficiency indicates whether there is sufficient supporting evidence to reach annotated verdicts; Retrieval refers to whether open-world evidence retrieval is performed; Unleaked represents whether annotated evidence contains temporal leakage such as including evidence published after claims.
Dataset Claim Evidence
Real Indep. Img. Suff. Retr. Unleaked
FEVER (Thorne et al., 2018) -
FEVEROUS (Aly et al., 2021) -
Liar-Plus (Alhindi et al., 2018)
Snopes (Hanselowski et al., 2019)
MultiFC (Augenstein et al., 2019)
AVeriTec (Schlichtkrull et al., 2023)
CLAIMDECOMP (Chen et al., 2024)
MOCHEG (Yao et al., 2023)
NewsCLIPpings (Luo et al., 2021) - - - -
InfoSurgeon (Fung et al., 2021) - - - -
Autosplice (Jia et al., 2023) - - - -
DGM (Shao et al., 2023) - - - -
MMFake (Liu et al., 2024b) - - - -
Verite (Papadopoulos et al., 2024) - - - -
COSMOS (Aneja et al., 2021) Mix - - - -
FACTIFY (Mishra et al., 2022) Mix - - - -
FACTIFY 2 (Suryavardan et al., 2023) Mix - - - -
MMOOC (Xu et al., 2024) Mix - - - -
Fauxtography (Zlatkova et al., 2019) - - - -
Fakeddit (Nakamura et al., 2020) - - - -
Qprop (Barrón-Cedeño et al., 2019) - - - -
FakeNewsNet (Shu et al., 2020) - - - -
MuMiN (Nielsen and McConville, 2022) - - - -
AVerImaTeC
@@ 4 Annotation Process
Claim Extraction & Normalization. Given a fact-checking article, an annotator extracts all claims from it, as multiple claims may pertain to a single event. The modality types of claims are annotated, and only image-text claims (i.e., textual claims paired with images) are forwarded to phase two. For identified image-text claims, sufficient context must be provided to ensure that the claim, specifically its textual component, can be understood independently of the surrounding article. The associated images are uploaded and normalized (e.g., separating collages or locating original, unaltered versions). Relevant metadata is also annotated. We exclude unverifiable claims (e.g., speculation or personal opinions) and claims where images are not used in verification or involve manipulated content.
@@ 5 Dataset Statistics
Data Distribution. We began with 2,353 fact-checking articles, and after discarding 210 inaccessible ones, we obtained 1,297 annotated image-text claims using the annotation pipeline described in the previous section. The splits of our dataset are temporally organized, and detailed statistics are presented in Table 2. We find that $23.7\%$ image-text claims include more than one claim image, highlighting the need to understand multiple visual inputs. On average, each claim is annotated with $2.92$ questions. Of these, $3.5\%$ questions have more than one answer, and $62.5\%$ are image-related, emphasizing the importance of visual context in verifying image-text claims. To address these questions, annotators frequently selected Image-search ($53.9\%$) as the answering method, indicating the necessity for tools that support image-centric information retrieval. Regarding answer types, $58.8\%$ are extractive, consistent with our annotation guidelines. Additionally, $1.6\%$ of answers are images themselves, underscoring the importance of supporting image retrieval as direct answers. A small proportion ($2.6\%$) of questions are marked as unanswerable, reflecting cases where no supporting evidence could be found online. Further metadata statistics, such as claim types and answer types, are provided in Appendix D.3. The dataset shows a label imbalance, with most claims being refuted, which is expected given that misleading content is more likely to be scrutinized.
@@ 6 Evaluation
Recent research (Akhtar et al., 2024) showed that reference-based evaluation of evidence with large language models (LLMs) aligns best with human assessments. We extend this framework to a multimodal setting where evidence from the web may comprise text and images. We transform each QA pair into an evidence statement following (Akhtar et al., 2024), where images are represented with special image tokens (e.g., [IMG_1]). For instance, the first QA pair in Figure 1 is transformed to [IMG_1] was published in 2016.”. We then conduct separate reference-based evaluations for the textual and visual components. For the textual part, we use a validation method similar to (Akhtar et al., 2024). If a matched evidence item is found within the ground-truth set, we proceed to a second step to compare the associated images. If image similarity falls below a threshold the match is deemed invalid due to image mismatch. We exploit Gemini-2.0-Flash (DeepMind, 2024) as the scoring model for both steps in the reference-based evaluation inspired by its power (Akhtar et al., 2024). We report evidence recall, defined as the percentage of ground-truth evidence instances successfully retrieved.
We performed alignment checks and robustness checks (against adversarial attacks) of different reference-based evaluation schemes (text-only, interleaved and the separated evaluation) in our setting (details in Appendix F.2 and F.3). Our separated reference-based evaluation achieved the highest alignment with human assessments, with a Spearman correlation coefficient ($\rho$(Spearman, 1904) of $0.332$ and a Pearson correlation coefficient ($r$(Pearson and Henrici, 1896) of $0.381$. Furthermore, the separated evaluation method is the most robust towards adversarial attacks.
@@ 7.1 Baselines
Question Generator. A straightforward approach is to generate all verification questions at once, given an image-text claim, as in prior work (Schlichtkrull et al., 2023). We refer to this strategy as paralleled question generation (PQG). However, since later questions often depend on earlier ones (e.g., as shown in Figure 1), we also propose a dynamic question generation (DQG) method, where each subsequent question is generated based on both the claim and the evolving evidence history. To combine the strengths of both approaches, we introduce a hybrid question generation (HQA) method, which generates the first few questions in parallel and then switches to dynamic generation for the remaining ones. All three strategies employ an MLLM for question generation. The prompts used for each strategy are detailed in Appendix I.3.
@@
Table 4: Baseline performance on verdict prediction and justification generation with ground-truth evidence (first block) and without accessing any external evidence (second block). NEE is for Not Enough Evidence and Conflict. is for Conflicting/Cherry-picking. Justi. is the for the performance of justification generation.
LLM MLLM Refuted Supported NEE Conflict. Overall Justi.
Gemini Gemini 0.84 0.92 0.52 0.00 0.82 0.50
@@ 8 Limitations
We made considerable efforts to prevent temporal leakage. Specifically, we provided annotators with custom search bars to retrieve online evidence published prior to the claim date. However, the exact dates of a small portion of claims ($5\%$) were unavailable. In such cases, we used the dates of the corresponding fact-checking articles, which may have been published a few days after the original claims. Additionally, we relied on Google Search and the Python package htmldate.find_date to estimate publication dates of web pages, which are coarse approximations. Furthermore, we exploited a reference-based evaluation strategy for both the generated questions and the retrieved evidence. Though it aligns well with human assessments, it has limitations in cases where model predictions are reasonable but not reflected in the reference annotations. In such cases, evaluation scores may be undeservedly low due to poor alignment with the references.

### v2

@@ 1 Introduction
We further introduce a baseline for image-text claim verification, which operates by generating evidence-seeking questions aimed at fact-checking and answering these questions with a set of expert tools. Since AVerImaTeC is the first image-text claim verification dataset to incorporate QA annotations that explicitly reflect reasoning paths and evidence, we develop a reference-based evaluation method to assess models’ generated questions and retrieved evidence. Using this evidence evaluation, we report conditional verdict accuracy which measures the correctness of predicted verdicts only when the associated evidence score exceeds a predefined threshold [^1] The dataset is available here: https://huggingface.co/datasets/Rui4416/AVerImaTeC. The code can be accessed here: https://github.com/abril4416/AVerImaTeC.
@@ 2 Related Works
Table 1: Comparison of fact-checking datasets. Indep. (independence) denotes whether extracted claims are context independent (e.g., understandable without fact-checking articles). Img., Suff., Retr., and Unleak. are abbreviations for image, sufficiency, retrieval and unleaked evidence, respectively. Sufficiency indicates whether there is sufficient supporting evidence to reach annotated verdicts; Retrieval refers to whether open-world evidence retrieval is performed; Unleaked represents whether annotated evidence contains temporal leakage such as including evidence published after claims.
Dataset Claim Evidence # Claims
Real Indep. Img. Suff. Retr. Unleak.
FEVER (Thorne et al., 2018) - 185,445
FEVEROUS (Aly et al., 2021) - 87,026
Liar-Plus (Alhindi et al., 2018) 12,836
Snopes (Hanselowski et al., 2019) 6,422
MultiFC (Augenstein et al., 2019) 36,534
AVeriTec (Schlichtkrull et al., 2023) 4,568
CLAIMDECOMP (Chen et al., 2024) 1,200
MOCHEG (Yao et al., 2023) 15,601
NewsCLIPpings (Luo et al., 2021) - - - - 988,283
InfoSurgeon (Fung et al., 2021) - - - - 30,000
Autosplice (Jia et al., 2023) - - - - 5,894
DGM (Shao et al., 2023) - - - - 230,000
MMFake (Liu et al., 2024b) - - - - 11,000
Verite (Papadopoulos et al., 2024) - - - - 1,000
COSMOS (Aneja et al., 2021) Mix - - - - 201,700
FACTIFY (Mishra et al., 2022) Mix - - - - 50,000
FACTIFY 2 (Suryavardan et al., 2023) Mix - - - - 50,000
MMOOC (Xu et al., 2024) Mix - - - - 364,000
Fauxtography (Zlatkova et al., 2019) - - - - 1,233
Fakeddit (Nakamura et al., 2020) - - - - 1,063,106
Qprop (Barrón-Cedeño et al., 2019) - - - - 51,294
FakeNewsNet (Shu et al., 2020) - - - - 23,196
MuMiN (Nielsen and McConville, 2022) - - - - 12,914
AVerImaTeC 1,297
@@
To capture the rationale in claim verification, a complex reasoning task, various reasoning representations have been explored, including natural logic (Strong et al., 2024), functional programs (Pan et al., 2023b), and QA (Chen et al., 2024; Qi et al., 2023; Pan et al., 2023a). In AVerImaTeC, we adopt QA as the reasoning representation, as the QA format is more intuitive and accessible for annotators, enabling efficient and consistent annotation by non-experts. Furthermore, other reasoning forms can often be mapped into QA-style representations, ensuring compatibility and flexibility for future extensions.
@@ 4 Annotation Process
Claim Extraction & Normalization. Given a fact-checking article, an annotator extracts all claims from it, as multiple claims may pertain to a single event. The modality types of claims are annotated, and only image-text claims (i.e., textual claims paired with images) are forwarded to phase two. For identified image-text claims, sufficient context must be provided to ensure that the claim, specifically its textual component, can be understood independently of the surrounding article. In some cases, metadata (e.g., date, location, speaker) is sufficient to disambiguate a claim. However, for ambiguities beyond metadata, such as unresolved coreference or missing referents, annotators are instructed to enrich the extracted claims so that they can be interpreted independently. The associated images are uploaded and normalized (e.g., separating collages or locating original, unaltered versions). Relevant metadata is also annotated. We exclude unverifiable claims (e.g., speculation or personal opinions) and claims where images are not used in verification or involve manipulated content.
@@ 5 Dataset Statistics
Data Distribution. We began with 2,353 fact-checking articles After discarding those that were inaccessible, did not focus on image-text claims, or contained unresolved annotation conflicts, we obtained 1,297 annotated image-text claims using the annotation pipeline described in the previous section. Further details on article sources and the filtering process are provided in Appendix D.1 and D.2, respectively. The splits of our dataset are temporally organized, and detailed statistics are presented in Table 2. We find that $23.7\%$ image-text claims include more than one claim image, highlighting the need to understand multiple visual inputs. On average, each claim is annotated with $2.92$ questions. Of these, $3.5\%$ questions have more than one answer, and $62.5\%$ are image-related, emphasizing the importance of visual context in verifying image-text claims. To address these questions, annotators frequently selected Image-search ($53.9\%$) as the answering method, indicating the necessity for tools that support image-centric information retrieval. Regarding answer types, $58.8\%$ are extractive, consistent with our annotation guidelines. Additionally, $1.6\%$ of answers are images themselves, underscoring the importance of supporting image retrieval as direct answers. A small proportion ($2.6\%$) of questions are marked as unanswerable, reflecting cases where no supporting evidence could be found online. Further metadata statistics, such as claim types and answer types, are provided in Appendix D.3. The dataset shows a label imbalance, with most claims being refuted, which is expected given that misleading content is more likely to be scrutinized.
@@ 6 Evaluation
Recent research (Akhtar et al., 2024) showed that reference-based evaluation of evidence with large language models (LLMs) aligns best with human assessments. We extend this framework to a multimodal setting where evidence from the web may comprise text and images. We transform each QA pair into an evidence statement following (Akhtar et al., 2024), where images are represented with special image tokens (e.g., [IMG_1]). For instance, the first QA pair in Figure 1 is transformed to “[IMG_1] was published in 2016.”. We then conduct separate reference-based evaluations for the textual and visual components. For the textual part, we use a validation method similar to (Akhtar et al., 2024). If a matched evidence item is found within the ground-truth set, we proceed to a second step to compare the associated images. If image similarity falls below a threshold the match is deemed invalid due to image mismatch. We exploit Gemini-2.0-Flash (DeepMind, 2024) as the scoring model for both steps in the reference-based evaluation inspired by its power (Akhtar et al., 2024). We report evidence recall, defined as the percentage of ground-truth evidence instances successfully retrieved.
We performed alignment checks and robustness checks (against adversarial attacks) of different reference-based evaluation schemes (text-only, interleaved and the separated evaluation) in our setting (details in Appendix F.2 and F.3). Our separated reference-based evaluation achieved the highest alignment with human assessments, with a Spearman correlation coefficient ($\rho$(Spearman, 1904) of $0.332$ and a Pearson correlation coefficient (rr(Pearson and Henrici, 1896) of $0.381$. Furthermore, the separated evaluation method is the most robust towards adversarial attacks.
@@ 7.1 Baselines
Question Generator. A straightforward approach is to generate all verification questions at once, given an image-text claim, as in prior work (Schlichtkrull et al., 2023). We refer to this strategy as paralleled question generation (PQG). However, since later questions often depend on earlier ones (e.g., as shown in Figure 1), we also propose a dynamic question generation (DQG) method, where each subsequent question is generated based on both the claim and the evolving evidence history. To combine the strengths of both approaches, we introduce a hybrid question generation (HQA) method, which generates the first few questions in parallel and then switches to dynamic generation for the remaining ones. All three strategies employ an MLLM for question generation, leveraging the model’s internal decomposition capability. The prompts used for each strategy are detailed in Appendix I.3.
@@
Table 4: Baseline performance on verdict prediction and justification generation with ground-truth evidence (first block) and without accessing any external evidence (second block). NEE is for Not Enough Evidence and Conflict. is for Conflicting/Cherry-picking. Justi. is for the performance of justification generation.
Evid. Source LLM MLLM Refuted Supported NEE Conflict. Overall Justi.
Ground-Truth Gemini Gemini 0.84 0.92 0.52 0.00 0.82 0.50
@@ 7.3 Analysis and Discussion
Although baselines used ground-truth evidence, they still achieved low justification generation scores. We attribute this to the limitations of ROUGE-1 for evaluating justification generation. To address this, we experimented with a reference-based evaluation method, Ev2R (Akhtar et al., 2024), which has been shown to align well with human assessments in open-ended generation tasks. Using this approach, the justification scores were much higher and more encouraging (details in Appendix G.2). Moving forward, we plan to explore and incorporate more appropriate evaluation methods for justification generation.
@@ 8 Limitations
AVerImaTeC has a relatively limited scale, as it is sourced from real-world claims and constructed through detailed human annotation. This is consistent with other human-annotated datasets of real-world claims, such as AVeriTeC (Schlichtkrull et al., 2023), which also contain only a few thousand claims. Since the claims in AVerImaTeC originate from fact-checking articles, the dataset may inherit biases inherent in these sources—for example, selection bias (Shin and Thorson, 2017; Barnoy and Reich, 2019), leading to imbalanced label distributions.
We made considerable efforts to prevent temporal leakage. Specifically, we provided annotators with custom search bars to retrieve online evidence published prior to the claim date. However, the exact dates of a small portion of claims ($5\%$) were unavailable. In such cases, we used the dates of the corresponding fact-checking articles, which may have been published a few days after the original claims. Additionally, we relied on Google Search and the Python package htmldate.find_date to estimate publication dates of web pages, which are coarse approximations.
Furthermore, we exploited a reference-based evaluation strategy for both the generated questions and the retrieved evidence. Though it aligns well with human assessments, it has limitations in cases where model predictions are reasonable but not reflected in the reference annotations. In such cases, evaluation scores may be undeservedly low due to poor alignment with the references.
9 Ethical Statement
The datasets and models described in this paper are not intended for use in truth-telling tasks, such as automated content moderation systems. The labels and justifications included in the dataset reflect only the evidence recovered by annotators and are therefore subject to the biases of both annotators and journalists. In addition, the QA annotations may introduce framing bias.
Annotators were instructed to prioritize evidence sources referenced in the original fact-checking articles, as we consider these sources, curated by professional fact-checkers, to be more trustworthy. Nonetheless, annotators were also permitted to draw on evidence retrieved by our customized search engine. Using the list of common misinformative sources from (Schlichtkrull et al., 2023), we observed that 8 answers relied on a flagged source. However, this list was not applied during annotation, as it may be incomplete or contain false positives. Moreover, source credibility is often context-dependent, varying across topics and over time, which makes a static list insufficient for reliable filtering.
We acknowledge that some fact-checking articles included in our dataset may have been exposed to large pre-trained models during their pre-training phase. This is an open and ongoing challenge in constructing human-annotated datasets. To help mitigate this issue, the splits of our dataset are temporally organized. In particular, if the training data of a language model is cut off prior to the temporal start of our test set, then data leakage from pre-training into evaluation cannot occur.
We did not anonymize the data in AVerImaTeC, as all claims are derived from publicly available journalistic sources and primarily concern public figures and events. Preserving these references is essential to ensure the accuracy and integrity of fact-checking. Nevertheless, we recognize that some individuals may not wish to appear in the dataset. Accordingly, we have established an opt-out policy: if any individual featured in the dataset, as a claim speaker, person depicted in an image, subject of a claim, or author of a fact-checking article underlying a claim, wishes to be removed, they may submit a request, and the relevant content will be deleted from the dataset.